OpenStudy (anonymous):

Choose the statement that correctly lists equivalent ratios. A. 1 : 6 = 20 : 120 = 5 : 30 B. 1 : 6 = 4 : 18 = 6 : 36 C. 1 : 8 = 20 : 120 = 3 : 24 D. 1 : 8 = 3 : 32 = 5 : 48

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Which ratio is equal to 3 : 5? A. 12 : 25 B. 12 : 20 C. 15 : 20 D. 12 : 15

OpenStudy (anonymous):

you want to find out which answer reduces to 3:5, so its best to think of them as fractions. for example:\[12:20\rightarrow \frac{12}{20}=\frac{\not4\cdot3}{\not4\cdot5}=\frac{3}{5}\rightarrow 3:5\] so your answer is 12: 20. Just think of reducing fractions. Likewise, you will see that the answer to your first question is really A.